In planning for your dream house, you are very particular with the probable amount of air particles. The bedroom has a measurement of 10 m by 7 m by 2 m. In the morning the temperature raises up to 23◦C and the air pressure is 98 mmHg. How much air particles may fit inside your broom?
1
Expert's answer
2020-11-09T14:11:36-0500

PV= nRT

P= 98mmHg = 13065.6Pa

V= 10×7×2 m3 = 140m3

T= 23°C = 273+23 = 296K

n=PV/RT

n = (13065.6×140)/(8.314×296)

n= 743.3 moles

1 mole = 6.02 ×1023 particles of air

743.3 moles = 743.3× 6.02×1023 particles of air

= 4.47×1026 particles of air